A labour market has 50,000 people in the labour force. Each month, a fraction p of employed


workers become unemployed (0 < p < 1) and a fraction q of unemployed workers become employed


(0 < q < 1).


(a) What is the steady-state unemployment rate?

(b) Under the steady-state, how many of the 50,000 in the labour force are employed and how many are employed each month? How many of the unemployed become employed each month?

(c) Suppose p = 0.08 and q = 0.32. What is the steady-state unemployment rate and how many workers move from employment to unemployment each month


1
Expert's answer
2021-12-22T13:56:10-0500

"Solution"

Let : "L=labor \ force ,\ e=employment, \ u=unemployment \\
L=e+u,\ so\ e=l-u\ ,u=l-e\\
\frac{u}{l}\ =unemployment \ rate,\frac{e}{l}=employment \ rate"

"S=rate \ of \ employed \ losing \ jobs\\
f=unemployed \ who \ get\ jobs"


"Steady \ state \ unemployment \ rate \ :\frac {u}{l}=\frac{s}{[s+f]}"


B)"\frac{u}{50000}=\frac{0<q<1}{(0<q<1+0<p<1)}"


C) "\frac{u}{50000}=\frac {0.32}{[0.08+0.32]}\\"

"0.4u=16000\\
u=4000"